# Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories

### v2

@@ Abstract
Numerical results and discussion.— While these states are low entropy zero-modes, they are degenerate with all the other zero-modes and are not easily identifiable in the spectrum. In the following, we introduce a potential that has these states as eigenstates. By analyzing the Shannon entropy and the bipartite entanglement entropy of the eigenstates of the Hamiltonian, we verify the existence of these low entropy states numerically. See the Supplementary Material for a brief overview of the calculation of these quantities as used in this work [53].
@@
From Eq. (7) we see that the sum of the height variables remains constant and equal to $(2i-S)$ over the entire sum. This does not mean that the sum over height variables is a valid potential because, in general, the height variables are not uniquely defined from the physical configuration. But they can be made unique for open boundary conditions, as described above. For example, in the case of a single-leg ladder, the sum of all height variables can be seen as the sum of the horizontal links of the top row. In general, for open boundary conditions, we write
@@
which correspond to choosing $i=1$ and $i=0$ in (7). We first study a $S=1$ single-leg ladder with $L$ plaquettes. This is a chain of plaquettes with periodic boundary conditions in the direction of the chain, but not the direction perpendicular to it. For such a ladder, there are exactly two distinct tilings of $2\times 1$ tiles. They correspond to placing the lower left corner of the tiles in either even or odd sites. These two states, coming from two different tilings with the same $i$, are not orthogonal but generate a two-dimensional scarred Hilbert space. We then expect four scars. Two degenerate low entropy states with energy $L\lambda/2$ from the tilings of state |$\ket{\psi_{+}}$ ($i=1$) and two with energy $-L\lambda/2$ from tilings of state |$\ket{\psi_{-}}$ ($i=0$). This is exactly what is observed when performing exact diagonalization on small systems, the results of which are presented in Fig. 3. The four low-entropy states are clearly identifiable and each of them is composed of the expected states.
@@
Figure 3: Left: Entanglement entropy of a $S=1$ ladder with $10\times 1$ plaquettes using the potential (8) with $\lambda=0.2$. Four mid-spectrum states with low entropy, shown in red, are visible. They are superpositions of the predicted states. For the remaining states, brighter colors indicate a higher density of states. Right: Entanglement entropy of $6\times 2$ $S=1$ plaquettes with periodic boundary conditions in the $x$ direction and $\lambda=0.2$ in the zero momentum sector. Many mid-spectrum states with low entropy are present. The towers marked in red contain the predicted states.
@@
For systems with a larger vertical direction, there are many more ways to tile the plane with $2\times 1$ and $1\times 2$ tiles. Adding a second leg is enough to make the number of possible tilings exponentially large, all of which will mix in the spectrum. Using the sum of the height variables as the potential, we expect towers of low-entropy states at energy ±$\pm L_{1}L_{2}\lambda/2$. These two towers are observed, as depicted in Fig. 3. There are many other low-entropy states at other energies, which make use of vertical tiles. They entail a further structure, that allows the number of height variables with value $1$ not to be the same as the ones with value 00. While these are still scars built up from zero-modes, we do not expect this construction to be generalizable beyond $2\times L$ systems and will not attempt a detailed characterization here. For periodic lattices, we studied the $4\times 4$ system with a generalized version of the potential. There the expected scar towers are also present. Details for this case can be found in the Supplementary Material [53].
Instead of the designed potential that was used to isolate the constructed scars, one might want to use the standard electric field term
@@
This potential exhibits scars that are constructed from zero-modes and allows for the identification of new tiling patterns. We identify three low entropy states in a $4\times 4$ lattice, which turn out to be constructed from zero-modes (see Fig. 4). Two of these are generalizable to arbitrary $L_{1}\times L_{2}$ lattices if both $L_{i}$ are even. We characterize one of them here, the remaining two scars are described in the Supplementary Material [53].
The scar has the form (Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories), just as the other scars we derived. However, the two plaquettes that make up tiles do not share a link, but neighbor each other diagonally in a pattern depicted in Fig. 4. This is still a zero-mode since each plaquette has one neighbor with height ±$\pm 1$ in all configurations and may therefore never take height $\mp 1$ itself. The state is an eigenstate of (10) because the number of neighbors with height ±$\pm 1$ is always the same for both plaquettes of each tile. This means that in all configurations, half of the links have values ±$\pm 1$ and the others are 00. Establishing the presence of these scars beyond $S=1$ is beyond the scope of this work. Systems with larger spins are less accessible through ED beyond single-leg ladders and we did not observe any of these scars for $S\leq 2$ in this case.
Refer to caption         Refer to caption
Figure 4: Left: Shannon entropy of $4\times 4$ $S=1$ plaquettes with periodic boundary conditions using the $E^{2}$ potential (10) with $g=0.2$ in the zero momentum and $+1$ charge conjugation sectors. Three mid-spectrum states with low entropy, are visible and colored in red. Bright colors indicate a higher density of states. Right: Illustration of a $S=1$ zero-mode that is an eigenstate of $\sum_{n}E_{n}^{2}$. A sublattice structure is built by singlet states of two plaquettes connected diagonally. Plaquettes connected by red or green lines form tiles.
@@
This demonstrates the existence of low entanglement entropy zero-modes in systems beyond $S=1$. For larger volume or higher spin systems, we have established theoretically the existence of such states in the earlier sections of this article.
Figure 5: Left: Entanglement entropy of a $S=2$ Ladder with 6 plaquettes with $\lambda=0.2$. Brighter colors indicate a higher density of states. Right: Shannon entropy of a $S=2$ Ladder with 6 plaquettes with $\lambda=0.2$. Four mid-spectrum states with low entropy are visible, colored in red. The remaining two predicted states are mixed with a tower of zero-modes at $E=0$.
Conclusions and outlook.— In this Letter, we demonstrate that QMBS exist across various 2D Abelian gauge theories. In the limit of zero coupling, a spectral symmetry emerges, leading to an exponentially large number of zero-modes. We show that it is possible to construct many states with area-law entanglement entropy, despite these states being in the middle of the spectrum. The number of these states grows exponentially with the volume, but they still comprise an exponentially small fraction of the Hilbert space, indicating weak ETH breaking.
We have analytically constructed QMBS for arbitrary truncated spins, revealing their existence beyond the reach of existing numerical methods, thereby marking a significant step towards understanding their role in the continuum limit, where the Hilbert space per link is not bounded. Our construction relies on the existence of numerous zero-modes, common in QLMs for arbitrary spin, though our method extends only to S=1 QLMs. This does not exclude that scars in higher spin QLMs exist but a different constructive approach might be needed. For $S=1$, some scars survive the presence of the standard $g\sum_{n}E^{2}_{n}$ potential. In a recent work [58], it has been argued that scars originally present on single-leg ladders $SU\left(2\right)$ gauge theories disappear once larger truncations of the Hilbert space are considered. Our results for $S=1$ show that these systems might be too simple to observe scars with the $E^{2}$ potential. An interesting generalization of our findings would involve extending the explicit QMBS construction to non-Abelian theories like $SU\left(2\right)$. From the experimental side, our theoretical description can guide future detection of QMBS and investigations of weak ergodicity breaking in systems inaccessible through classical simulations.

### v3

@@ IV Numerical results and discussion
While the derived scars are low entropy zero-modes, they are degenerate with all the other zero-modes and are not easily identifiable in the spectrum. In the following, we introduce potentials that have these states as eigenstates. By analyzing the Shannon entropy and the bipartite entanglement entropy of the eigenstates of the Hamiltonian, we verify the existence of these low entropy states numerically. Concretely, we calculate the Shannon entropy with respect to the electric field basis {|$\{\ket{\phi_{i}}\}$. For a state |$\ket{\psi}=\sum_{i}c_{i}\ket{\phi_{i}}$, the Shannon entropy has the value
$S=-\sum_{i}|c_{i}|^{2}\log|c_{i}|^{2}.$ (14)
For the entanglement entropy, we split the system into two subsystems $A$ and $B$ of equal size. The bipartite entanglement entropy is then given by,
$S_{\frac{L}{2}}=-\text{Tr}\rho_{A}\log\rho_{A}=-\text{Tr}\rho_{B}\log\rho_{B}.$ (15)
Explicitly, we always partition the system by dividing it vertically into two halves. We consider the bordering vertical links on the left as belonging to the subsystem while the ones on the right do not. See e.g. the Supplementary Material of [30] for details on how to calculate the entanglement entropy in these systems.
@@
We start by isolating the scars of the form of (24). We observe that the sum of the height variables remains constant and equal to $(2i-S)$ for each term. This does not mean that the sum over height variables is a valid potential because, in general, the height variables are not uniquely defined from the physical configuration. However, they can be made unique for open boundary conditions, as described above. For example, in the case of a single-leg ladder, the sum of all height variables can be seen as the sum of the horizontal links of the top row. In general, for open boundary conditions, we write
@@
which correspond to choosing $i=1$ and $i=0$ in (11). We first study a $S=1$ ladder with $L\times 1$ plaquettes. This is a chain of plaquettes with periodic boundary conditions in the direction of the chain, but not the direction perpendicular to it. For such a ladder, there are exactly two distinct tilings of $2\times 1$ tiles. They correspond to placing the lower left corner of the tiles in either even or odd sites. These two states, coming from two different tilings with the same $i$, are not orthogonal but generate a two-dimensional scarred Hilbert space. We then expect four scars. Two degenerate low entropy states with energy $L\lambda/2$ from the tilings of state |$\ket{\psi_{+}}$ ($i=1$) and two with energy $-L\lambda/2$ from tilings of state |$\ket{\psi_{-}}$ ($i=0$). This is exactly what is observed when performing exact diagonalization on small systems, the results of which are presented in Fig. 5. The four low-entropy states are clearly identifiable and each of them is composed of the expected states, as shown in the Supplementary Material [56] by looking at the amplitude of the scars in the electric field basis.
@@
Figure 5: Entanglement entropy for two different volumes with periodic boundary conditions in the $x$ direction, $S=1$, and height potential (16) with $\lambda=0.2$. Brighter colors indicate a higher density of states. Left: $10\times 1$ ladder. Four mid-spectrum states with low entropy, shown in red, are visible. They are superpositions of the predicted states. Right: $6\times 2$ volume in the zero momentum sector. Many mid-spectrum states with low entropy are present. The towers marked in red contain the predicted states.
@@
For systems with a larger vertical direction, there are many more ways to tile the plane with $2\times 1$ and $1\times 2$ tiles. Adding a second leg is enough to make the number of possible tilings exponentially large, all of which will mix in the spectrum. Using the sum of the height variables as the potential, we expect towers of low-entropy states at energies ±$\pm L_{1}L_{2}\lambda/2$. These two towers are observed, as depicted in Fig. 5. There are many other low-entropy states at other energies. While we observed these states to be built from zero-mode tiles, we do not expect their construction to be generalizable beyond $2\times L$ systems and will not attempt a detailed characterization here.
All scars we have constructed also exist in periodic boundary conditions. Using the sum of the height variables as the potential is not well defined, as previously discussed. Instead, for the scars of the form (24), we use the minimum number of kinetic operators that must be applied to the state |$\ket{\textbf{0}}$ to arrive at any given electric field basis state |$\ket{\phi_{i}}$. This is similar to the Hamming distance as used in e.g. [64] generalized for higher spin. We therefore call this the Hamming distance potential. The entropy using this potential is shown in the left panel of Fig. 6, where the tower of predicted scars is visible. Additional towers are present at non-integer multiples of $\lambda$, which are not characterized here.
For the other types of scars described for $S=1$, all contributing states in (12) and (13) have the same number of links with electric field zero. In the first case, half of the links are zero while in the second no link is zero. These scars can then be isolated by the standard electric field term
@@
The third scar, described in the Supplemental Material [56], is also an eigenstate of this potential. These are precisely the three scars observed in the right panel of Fig. 6.
Figure 6: Shannon entropy for a $4\times 4$ system with periodic boundary conditions in both directions, $S=1$, and different potentials. Shown states are in the zero momentum and $+1$ charge conjugation sectors. Bright colors indicate a higher density of states. Left: Hamming distance potential with $\lambda=0.2$. Many mid-spectrum states with low entropy are visible. The tower colored in red at $E=1.6$ contains the predicted scars. Right: Standard electric field term (18) with $g=0.2$. Three mid-spectrum states with low entropy are visible and colored in red.
@@
This numerically demonstrates the existence of low entanglement entropy zero-modes in systems beyond $S=1$. For larger volume or higher spin systems, we have established the existence of such states analytically in the earlier sections of this article.
Figure 7: Entropies for a $S=2$ ladder with 6 plaquettes and $\lambda=0.2$. Brighter colors indicate a higher density of states. Left: Entanglement entropy. Right: Shannon entropy. Four mid-spectrum states with low entropy are visible, colored in red. The remaining two predicted states are mixed with a tower of zero-modes at $E=0$.
V Conclusions and outlook
In this work, we demonstrate that Quantum Many-Body Scars (QMBS) exist across numerous 2D Abelian pure gauge theories. In the limit of zero coupling, a spectral symmetry emerges, leading to an exponentially large number of zero-modes. We show that it is possible to construct many states with area-law entanglement entropy within this subspace. Since these are mid-spectrum states of a non-integrable model, they constitute QMBS. The number of scars we could identify grows exponentially with the volume, but comprise an exponentially small fraction of the Hilbert space. This indicates weak breaking of the Eigenstate Thermalization Hypothesis.
Through the analytical construction of the scars, we can strategically add a potential to the Hamiltonian that isolates them from other eigenstates. This allows for the numerical verification of their presence for moderate volumes and values of the spin. Different choices of the potential can isolate different scars. If the new potential does not have scar states as eigenstates, they will disappear from the spectrum. An interesting question for future work is whether this type of analytical construction can be generalized in the presence of bosonic or fermionic matter.
For $S=1$, some scars survive the presence of the standard $g\sum_{n}E^{2}_{n}$ potential, if the lattice size is even in both directions and under periodic boundary conditions. It would be interesting to understand if a similar dependence on the geometry could occur in other gauge theories, such as $SU\left(2\right)$, where scars were not found for larger truncations in single-leg ladders [65].
Our analytical construction of QMBS extends to arbitrary truncated spins and arbitrarily large volumes. This result establishes the presence of scars well beyond the reach of existing numerical methods. It can be used to guide future quantum simulations of gauge theories, in a regime beyond classical simulations, towards interesting non-equilibrium phenomena. From the theoretical point of view, it constitutes a step towards understanding the role of scars when taking the continuum limit, where the Hilbert space per link is not bounded.
@@ V Conclusions and outlook
Supplementary Material
I Symmetries of the systems
The symmetries of the Hamiltonians considered will depend on the choice of the potential. We will always have
  • Translation Symmetry along the horizontal axis $E_{ni}\rightarrow E_{n+\hat{1}i}$, $U_{n+\hat{1}i}\rightarrow U_{n+\hat{1}i}$. Translation along the vertical axis is not always present due to boundary conditions;
  • Reflection Symmetries with respect to the two axis ${\cal I}_{x}$ and ${\cal I}_{y}$. Explicitly, ${\cal I}_{x}$ is given by $E_{\left(n_{1},n_{2}\right)1}\rightarrow E_{\left(n_{1},-n_{2}\right)1}$, $E_{\left(n_{1},n_{2}\right)2}\rightarrow-E_{\left(n_{1},-n_{2}\right)2}$ $U_{\left(n_{1},n_{2}\right)1}\rightarrow U^{\dagger}_{\left(n_{1},-n_{2}\right)1}$, $U_{\left(n_{1},n_{2}\right)2}\rightarrow U^{\dagger}_{\left(n_{1},-n_{2}\right)2}$ and analogous for ${\cal I}_{y}$. These symmetries are important to prove the existence of an exponential number of zero-modes;
  • Charge Conjugation, which is characterized by the field transformations $E_{ni}\rightarrow-E_{ni}$ and $U_{ni}\rightarrow U^{\dagger}_{ni}$.
When explicitly referenced, we restricted our exact diagonalization to the zero momentum sector and to the +1 charge conjugation sector. Otherwise, the plots are in the electric field basis and no symmetries beyond applying Gauss’ law and being restricted to the zero winding sector are used.
II Index Theorem and an Exponential Number of zero-modes
In [62] it was shown that non-integrable models can still exhibit an exponential number of eigenstates with zero energy. We review this result and adapt it to our models of interest. For this section, we will refer to plaquette flipping terms as the kinetic part of the Hamiltonian
$K=\sum_{n}\left(U_{n1}^{\dagger}U_{n+\hat{1}2}^{\dagger}U_{n2}U_{n+\hat{2}1}+\mathrm{h.c.}\right).$ (19)
We only need to assume that $U_{ni}/U_{ni}^{\dagger}$ are operators that raise/lower the value of the electric field on the link by 1. This makes the result very robust with respect to a wide range of formulations.
We define a set of link operators $\zeta_{ni}$, which are diagonal in the electric field basis. For each link we have
$\zeta\ket{\varepsilon}=\left(-1\right)^{\varepsilon}\ket{\varepsilon}.$ (20)
We further construct
${\cal C}=\prod_{n=0}^{L_{1}-1}\prod_{a=0}^{L_{2}/2-1}\zeta_{\left(n,2a\right)1},$ (21)
where we have assumed $L_{1}$ to be even. We can see that ${\cal C}$ anti-commutes with the kinetic Hamiltonian $\left\{{\cal C},K\right\}=0$. This implies that if |$\ket{E}$ is an eigenstate with energy $E$, then ${\cal C}\ket{E}$ has energy $-E$. In the subspace generated by eigenstates with zero energy, we can diagonalize ${\cal C}$ and $K$ together.
Let now ${\cal I}_{y}$ be the unitary transformation that reflects along the vertical axes, as defined above. This is a symmetry of the Hamiltonian and also commutes with ${\cal C}$. We can now use the existence of ${\cal I}_{y}$ and ${\cal C}$ to show the existence of an exponential number of zero-modes.
First, note that the number of zero-modes $N_{0}$, is bounded from below by $N_{0}\geq\left|\mathrm{tr}\left({\cal CI}_{y}\right)\right|$. We can see this by adopting a basis that diagonalizes both $K$ and ${\cal I}_{y}$. Such eigenstates are always mapped to a different energy eigenstate by ${\cal C}$ unless the energy is zero. As a consequence, only zero-modes contribute to the trace. Each eigenstate, on that sector, will contribute with a value in the interval $\left[-1,1\right]$. This shows that the number of zero-modes can never be smaller than $\left|\mathrm{tr}\left({\cal CI}_{y}\right)\right|$.
We can now show that the trace is actually exponential on the volume by considering the electric field basis. These are all eigenstates of ${\cal C}$, but the only ones surviving the trace are the ones that respect a reflection symmetry along the vertical axis. Such states are fixed by determining the state in half of the volume. The number of these states corresponds to the square root of the total number of states in the whole Hilbert space. By restricting ourselves to specific sectors of the Hilbert space, the two halves are not independent. Nonetheless, we still expect the number of zero-modes to remain exponentially large on the volume.
III Low Entropy Zero-Modes in Truncated Link Models
Here we describe a generic mechanism to obtain low-entropy mid-spectrum states in the TLM, from the vast space of zero-modes. We start by searching for zero-modes |$\ket{\psi_{z}}$, with a generic decomposition following |$\ket{\psi_{z}}=\sum_{n}c_{i}\ket{\phi_{i}}$, where {|$\left\{\ket{\phi_{i}}\right\}$ is the electric field basis. By definition, a zero-mode follows
$K\ket{\psi_{z}}=\sum_{i,n,\sigma}c_{i}H^{\sigma}_{n}\ket{\phi_{i}}=0,$ (22)
where $\sigma\in\{+,-\}$. This is very general and must be satisfied by all zero-modes.
Spin-$S$ TLMs have an especially simple structure, which we can use to find low-entropy zero-modes. If $H^{\sigma}_{n}\ket{\phi_{i}}\neq 0$ then $||H^{\sigma}_{n}\ket{\phi_{i}}||=1$ for all $n$, $i$ and $\sigma$. This suggests that we can look for solutions where each term in (22) is canceled by exactly one other term in the sum. In other words, for each non-null term $c_{i}H^{\sigma}_{n}\ket{\phi_{i}}$, there exists a term $c_{k}H^{\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}$ with $c_{k}=-c_{i}$ such that $H^{\sigma}_{n}\ket{\phi_{n}}=H^{\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}$ for some $\sigma^{\prime},n^{\prime}$. This introduces the minimal possible number of non-zero weights to cancel a given $H^{\sigma}_{n}\ket{\phi_{i}}$. In TLMs, states |$\ket{\phi_{i}}$, |$\ket{\phi_{k}}$ that obey $H^{\sigma}_{n}\ket{\phi_{i}}=H^{\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}\neq 0$, are related by |$\ket{\phi_{k}}=H^{-\sigma^{\prime}}_{n^{\prime}}H^{\sigma}_{n}\ket{\phi_{i}}=H^{\sigma}_{n}H^{-\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{i}}$.
We then focus on two options to cancel a term $c_{i}H^{\sigma}_{n}\ket{\phi_{i}}$ in Eq. (22) for TLMs:
  1. 1.
    |$\ket{\phi_{i}}$ is already annihilated by $H^{\sigma}_{n}$;
  2. 2.
    The zero-mode also includes the term $-c_{i}\ket{\phi_{k}}=-c_{i}H^{-\sigma^{\prime}}_{n^{\prime}}H^{\sigma}_{n}\ket{\phi_{i}}$ with some $\sigma^{\prime}$, $n^{\prime}$.
The second option not only assures that $H^{\sigma}_{n}\ket{\phi_{i}}-H^{\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}=0$, but also that $H^{-\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{i}}-H^{-\sigma}_{n}\ket{\phi_{k}}=0$. This means that the contributions $c_{i}H^{-\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{i}}$ and $c_{k}H^{-\sigma}_{n}\ket{\phi_{k}}$ also cancel each other. What remains are the terms $c_{i}H^{-\sigma}_{n}\ket{\phi_{i}}$, $c_{i}H^{\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{i}}$, $c_{k}H^{-\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}$ and $c_{k}H^{\sigma}_{n}\ket{\phi_{k}}$ as well as all terms that act on different sites. These also need to be canceled to get a zero-mode. If it is possible to find a set of states that manage to cancel all contributions in this fashion, this state is a good candidate to be a zero-mode with anomalously low entropy.
Canceling the remaining contributions, could, in principle, be achieved by considering any of these states and acting upon them with another kinetic operator. For example, $H^{\sigma}_{n}\ket{\phi_{k}}$ could be cancelled by adding a state of the form $H^{\sigma^{\prime\prime}}_{n^{\prime\prime}}H^{\sigma}_{n}\ket{\phi_{k}}$, for arbitrary choices of $n^{\prime\prime}$ and $\sigma^{\prime\prime}$ that do not annihilate the state. To arrive at the zero-modes in arbitrary integer spin, as described in the main text, we choose to cancel them using the same operators as before, so with the state |$\ket{\phi_{l}}=H^{-\sigma^{\prime}}_{n^{\prime}}H^{\sigma}_{n}\ket{\phi_{k}}$. This cancels two contributions at the same time, $H^{\sigma}_{n}\ket{\phi_{k}}$ and $H^{-\sigma^{\prime}}_{n^{\prime}}\ket{\phi_{k}}$, further reducing the number of states that need to be included. We then build chains of such states, where both ends are annihilated by the remaining kinetic operators. This way, we build the $2\times 1$ zero-mode state that can be tiled according to the product defined in the following section.
All analytically constructed zero-mode scars that have been derived for the $S=1/2$ QLM [31, 32] have this structure and therefore an equal magnitude of weights for all contributing electric field basis states.
The outlined strategy can be used to construct more zero-mode scars in systems that are not accessible through ED. For other formulations, like QLMs, it will still be true that states that cancel each other in (22) differ by two applications of a kinetic operator. This insight was crucial to understanding the low entropy states observed here for TLMs and may prove useful in the analysis of other systems.
IV The tiling product
There are multiple options for generalizing scars found in small systems to arbitrarily large lattices. The most straightforward approach is to apply tensor products in the electric field basis. This is only possible if the links at the boundary agree and the resulting system still follows Gauss’ law. We generalize this to the tiling product $\odot$, which acts like a tensor product on the dual representation instead. It is defined as
|$\ket{M}\odot\ket{M^{\prime}}=\ket{MM^{\prime}},$ (23)
i.e. the matrices of height variables are concatenated. This is not the tensor product, since the values of the boundary links are added. If the concatenation of $M$ and $M^{\prime}$ would give rise to an invalid gauge field configuration, i.e. some neighboring height variables satisfy $|h_{i}-h_{j}|>S$, we will say that |$\ket{M}$ and |$\ket{M^{\prime}}$ do not tile and define their tilling product to give the null state. More generally, we are interested in tiling products of the form $\left(\sum_{a}c_{a}\ket{M_{a}}\right)\odot\left(\sum_{b}c_{b}\ket{M^{\prime}_{b}}\right)$. We set this product to result in the null state if any state in the left-hand side does not tile with any state in the right side, otherwise, we just apply the distributive law. In other words, the tiling is distributive if all the states in one entry tile with all the states in the other.
For example, states that only contain height variables $h_{i}$ with $\max_{i}(h_{i})-\min_{i}(h_{i})\leq S$, will always tile with themselves. If the difference is larger, one might need to be careful which plaquettes can touch, to prevent neighboring height variables from differing by more than $S$.
This tiling product gives non-null states (specifically the scars) for the tiles constructed in the main text since $\max_{i}(h_{i})-\min_{i}(h_{i})=S$. In general, scars do not need to have a tiling structure. One eigenstate of $E^{2}$ is not composed of tilings. It is described in the following section.
V A $4\times 4$ Scar for the $E^{2}$ potential
The remaining observed scar referred to in the main text seems exclusive of the $4\times 4$ system. While the fundamental tiling structure still applies, it comes from a different tiling of the states described in the main text. The two plaquettes in a tile lie in the same row or column but have a distance of 2. Explicitly
|$\ket{\psi_{s}}=\frac{1}{16}\prod_{(n,n^{\prime})\in T}\left(H^{+}_{n}-H^{+}_{n^{\prime}}\right)\ket{\mathbf{0}},$ (24)
where the set of tiles $T=\left\{T_{1},\dots,T_{8}\right\}$ is presented in Fig. 8.
Figure 8: Schematic representation of a tiling that produces a scar for the $E^{2}$ potential in a $4\times 4$ lattice.
This tiling guarantees that each $T_{i}$ shares the same number of neighboring height variables of value 0 and 1 and therefore will be an eigenstate of the potential with eigenvalue $16g$.
VI Amplitudes of Scars
The amplitudes of the isolated low-entropy states match the predictions of the constructed scars. We demonstrate this in the case of the negative energy scars in $S=1$ single-leg ladders.
The predicted scars with negative energy are generated by tilings of |$\ket{\psi_{-}}$ defined in the main text. In the single-leg ladder, only two tilings exist. This generates two states of equal energy, which we call |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$. They are linearly independent and will be found in a superposition |$\ket{\psi_{S}}=a\ket{\psi_{1}}+b\ket{\psi_{2}}$. All non-zero amplitudes $\braket{\phi_{i}}{\psi_{1,2}}$, in the electric field basis, are ±$\pm\frac{1}{\sqrt{2^{L/2}}}$.
Figure 9: Overlap of electric field basis states |$\ket{\phi_{i}}$ with the lowest entropy scar in Fig. 3 of the main text. It depicts the entanglement entropy of $10\times 1$ $S=1$ plaquettes with periodic boundary conditions in the $x$ direction. Most states have zero overlap.
There are therefore four different possible magnitudes of the amplitudes $A=\braket{\phi_{i}}{\psi_{S}}$:
  1. 1.
    The basis state |$\ket{\phi_{i}}$ has zero overlap with both |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ and therefore the amplitude is zero.
  2. 2.
    The basis state has non-zero overlap with only |$\ket{\psi_{1}}$ and has the amplitude $A=\pm\frac{a}{\sqrt{2^{L/2}}}$.
  3. 3.
    The basis state has non-zero overlap with only |$\ket{\psi_{2}}$ and has the amplitude $A=\pm\frac{b}{\sqrt{2^{L/2}}}$.
  4. 4.
    The basis state has non-zero overlap with both |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ and has the amplitude $A=\pm\frac{a+b}{\sqrt{2^{L/2}}}$.
The fourth case only occurs for the states |$\ket{10101010...}$ and |$\ket{01010101...}$, which exist in both tilings. The eight different amplitudes are visible in Fig. 9. It shows the overlap of the electric field basis states |$\ket{\phi_{i}}$ with a scar of a $10\times 1$ ladder, with spin $S=1$, under periodic boundary conditions in the $x$ direction.